OpenStudy (anonymous):

Solve. h minus two and three-fourths equals one and one-half h equals one-half h equals three-fourths h equals three and one-half h equals four and one-fourth

OpenStudy (anonymous):

@Glizzy_Nedd

OpenStudy (rational):

\(\large h - 2 \frac{3}{4} = 1\frac{1}{2}\)

OpenStudy (rational):

like that ?